On January 1 2018, Poultry Processing Company purchased a freezer and related installation equipment for $65,400. The equipment had a three-year estimated life with a $4,200 salvage value. Straight-line depreciation was used. At the beginning of 2020, Poultry Processing revised the expected life of the asset to four years rather than three years. The salvage value was revised to $3200.

Required

Compute the depreciation expense for each of the four year, 2018-2021

Depreciation Expense

2018

2019

2020

2021
Answer #1

Depreciation (2018) = (65400-4200) / 3 = $20,400

Depreciation (2019) = (65400-4200) / 3 = $20,400

Net book value at the end of 2019 = 65400 - 20400*2 = $24,600

Depreciation (2020) = (24600-3200)*1/2 = $10,700

Depreciation (2021) = (24600-3200)*1/2 = $10,700